OpenStudy (anonymous):

how can I express the volume of a solid rectangular box as a function of it's width. height = x, width = 2x, lenght = 3x

OpenStudy (aroub):

l*w*h

OpenStudy (anonymous):

\[V(w) = l*w*h = w/2 * w * w * (3/2) = (3/4) w ^{3}\]